OpenStudy (anonymous):

**Medal to everyone who answers and explains well Directions: Find the coordinates of the midpoint of each segment. 1. line QR with endpoints Q(0,5) and R(6,7) 2.line JK with endpoints J(1,-4) and K(9,3) Please show work :) Geometry just isn't my thing...

OpenStudy (anonymous):

Use the formula( (x1+x2)/2,(y1+y2/2) )

OpenStudy (anonymous):

\[midpoint = x _{1}+_{2}/2 , y_{1}+y _{2}/2\] 0+6/2, 5+7/2 =3,6 is midpoint

OpenStudy (anonymous):

thanks everyone!

OpenStudy (anonymous):

same for no.2 1+9/2 = 10/2 =5 -4+3/2 = -1/2 so midpoint is 5,-1/2